OpenStudy (anonymous):

j-mart was having a major sale. if susan received two successive discounts of 10 % and 15% on an item that cost $90.00 what is the final discounted cost? j-mart was having a major sale. if susan received two successive discounts of 10 % and 15% on an item that cost $90.00 what is the final discounted cost? @Mathematics

OpenStudy (a_clan):

after 10% discount, (90 - 0.1*90)=81 after 15% discount, 81- 0.15*81= 68.85

OpenStudy (anonymous):

if the successive discounts were 15 % and 10 % what would the $90.00 item coost?

OpenStudy (a_clan):

68.85 still

OpenStudy (anonymous):

if the storegave one discount of 25% instade of two successive discounts, how much would that $90.00 item cost?

OpenStudy (a_clan):

67.5

OpenStudy (anonymous):

show work please

OpenStudy (a_clan):

90 - (25/100)*90 = 67.5

OpenStudy (anonymous):

thanks so muchhhh

OpenStudy (a_clan):

np